40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jul 10 02:15:36 2014, Jenne, Update, LSC, QPD dark noise checkout 
    Reply  Fri Jul 11 20:05:34 2014, Jenne, Update, LSC, QPD dark noise checkout TransQPDY_darkNoise_vs_TransimpedanceGain_9July2014.pdfTransQPDY_darkNoise_vs_WhiteningGain_9July2014.pdfsqrtTrans_sensitivity.pngsqrtTrans_Noise_vs_CARMoffset.png
       Reply  Tue Jul 15 21:43:28 2014, Jenne, Update, LSC, QPDs back 
Message ID: 10206     Entry time: Tue Jul 15 21:43:28 2014     In reply to: 10187
Author: Jenne 
Type: Update 
Category: LSC 
Subject: QPDs back 

I removed the dumps in front of the trans QPDs.  The Yend QPD needed re-normalization, so I did that.

ELOG V3.1.3-